OpenStudy (anonymous):

Solve: 2x + 10 + 26 = 212

OpenStudy (anonymous):

2x + 10 + 26 = 212 First, add like terms: 2x + 36 = 212 Then subtract: 2x = 176 Then divide by 2: x=88

OpenStudy (uri):

2x=212-10-26 2x=176 x=176/2 divide and ans

OpenStudy (anonymous):

use wolfram alpha go on Google and type it wolfram alpha it should help